OpenStudy (anonymous):

It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

Directrix (directrix):

I'm looking at this approach: http://www.purplemath.com/modules/workprob.htm

OpenStudy (anonymous):

Sorry, LaTex/sigma Equation is not working at this time. 1/x + 1/(x + 15) = 1/4 1 car per x hours + 1 car per (x+15) hours equals 1 car in 4 hours. Solve for x. x = 5 Brian takes 5 + 15 = 20 hours